A bottle opener requires that a force of 25N must be applied to the handle in other to lift a bottle cap 0.90cm the bottle cap has a VR of 800 and an effeciency of 75%. What force is applied to the bottle cap?. How far does the handle of the opener moves?


1
Expert's answer
2021-02-22T10:20:27-0500

Given quantities: "F = 25N" "h = 0.9cm = 9*10^{-3}m" "\eta=0.75"

VR - vertical radius "R = 0.8cm"

"\eta = \large\frac{A_{useful}}{A_{total}}" "= \large\frac{F*h}{F_2*R}" "\to F_2 = \large\frac{Fh}{R\eta}=\frac{25*9*10^{-3}}{8*10^{-3}*0.75}" = 37.5N